Two red onion plants are crossed to give offspring yielding onions are: red,122 : yellow, 29: white, 9.

What is the most appropriate null hypothesis for these data?

A. 9:3:3:1, B. 9:7, C. 12:3:1

What is your best estimate of the chi-square statistic?

A. 0, B. 0.103, C. 0.0167

What are the degrees of freedom?

A. 1, B. 2, C. 3, D. 4, E. none
Answer #1

Null hypothesis: C. 12:3:1

Since there are three phenotypes in the progeny, this seems to be a case of dominant epistasis where all A-/B- and aa/B- progeny show the same phenotype (numbered 12, red). The A-/bb phenotype numbers 3 (yellow) and the aabb phenotype numbers 1 (white).

The best estimate of the chi square statistic is 0.103

Total progeny = 122+29+9 = 160. From the expected ratio 12:3:1, expected progeny are 160x 12/16 = 120 red, 160x 3/16 = 30 yellow and 160x 1/16 = 10 white.

Degree of freedom = 1 (since there are two genes involved, dof becomes 2-1= 1)